OpenStudy (anonymous):

If 2 over 3 ton of gravel covers 3 over 4 of a driveway, how many tons of gravel are required to cover the entire driveway? 1 over 3 ton 1 over 2 ton 8 over 9 ton 11 over 12 ton

OpenStudy (anonymous):

i will fan and medal

OpenStudy (anonymous):

help plz

OpenStudy (anonymous):

anyone

OpenStudy (anonymous):

i will fan and medal if anyone helps

OpenStudy (anonymous):

can someone help plz

jimthompson5910 (jim_thompson5910):

Form a ratio and multiply both parts of the ratio by 4/3 to turn the "3/4" into "1" (to represent 1 full driveway) \[\Large \frac{2}{3} \ \text{ton} \ : \frac{3}{4} \ \text{driveway}\] \[\Large \color{red}{\frac{4}{3}*}\frac{2}{3} \ \text{ton} \ : \color{red}{\frac{4}{3}*}\frac{3}{4} \ \text{driveway}\] \[\Large \frac{4*2}{3*3} \ \text{ton} \ : \frac{4*3}{3*4} \ \text{driveway}\] \[\Large \frac{8}{9} \ \text{ton} \ : \frac{12}{12} \ \text{driveway}\] \[\Large \frac{8}{9} \ \text{ton} \ : 1 \ \text{driveway}\] So we need \(\Large \frac{8}{9}\) of a ton of gravel.

OpenStudy (anonymous):

thx